OpenStudy (anonymous):

Need help with this 1 question..thanks, medals if right or if not right ;) Solve: (x - 8)(x + 11) = 0.

OpenStudy (callisto):

To solve it, you need to know either (x - 8) =0 or (x + 11) = 0 , as something multiplying 0 give 0 So (x - 8)=0 or (x + 11) = 0 x = 8 or x=-11 Do you understand it?

OpenStudy (anonymous):

Oh okay. Haha duh..

OpenStudy (anonymous):

x=3.14 or x=2.718 ?

OpenStudy (anonymous):

woohoo!

OpenStudy (anonymous):

i dont think thats right... read the person above..it makes sense lol

OpenStudy (anonymous):

callisto's right...

OpenStudy (anonymous):

i was just testing the validity of what the question says...

OpenStudy (anonymous):

oh so callisto was wrong.. i put no solution and got the question wrong... i dont understtand it now!

OpenStudy (anonymous):

no. callistos right...

OpenStudy (anonymous):

:{

OpenStudy (callisto):

Nah? Was I wrong?

OpenStudy (anonymous):

Ya lol

OpenStudy (anonymous):

the callisto's way of doing this is correct.

OpenStudy (callisto):

The solutions are x = 8 or x=-11 Are you sure?

OpenStudy (anonymous):

yesss Callisto is right...

OpenStudy (callisto):

* are you sure I was wrong?

OpenStudy (accessdenied):

AB = 0 <=> A = 0 or B = 0 sounds about right to me

OpenStudy (anonymous):

OH. you were right all along. i read it wrong. i thought the anwser was 0

OpenStudy (anonymous):

the owl with the black hat is wrong....

OpenStudy (anonymous):

Im sorry! Lol. dee D freaking dee

OpenStudy (anonymous):

haha thats you? ............